THE UNIVERSITY OF CHICAGO 2007 Summer VIGRE Program for Undergraduates

Preference Form

0. Please give your name and say whether you are in the apprentice or full program.

1. Please list the numbers of the courses that you feel seem most likely to interest you, in order of preference. (If you are in the apprentice program, please list 5 first.)

2. Please list any other topics that you feel you would most like to discuss most with a graduate student mentor.

LIST OF COURSES

1. LASZLO BABAI. DISCRETE MATHEMATICS. Eight weeks. First four weeks: Numbers, groups, and graphs. Second four weeks: Finite and transfinite combinatorics: the world of Paul Erdős. 2a. GREGORY LAWLER. RANDOM WALKS AND BROWNIAN MOTION. Weeks one and two. 2b. ROBERT FEFFERMAN. AN INTRODUCTION TO THE CONVERGENCE OF FOURIER SERIES. Weeks three and four. 2c. THOMAS FIORE: MATHEMATICS AND MUSIC. Weeks five and six. 3a. BENSON FARB. SOME GEOMETRY AND TOPOLOGY. Weeks one through four. 3b. URI BADER. A TOPIC IN GEOMETRY. Weeks five and six. 3c. MATTHEW KERR. A TOPIC IN ALGEBRAIC GEOMETRY. Weeks seven and eight.

4a. PETER MAY. SURFACES, ALGEBRAS, AND TQFT’s. Weeks one through four. 4b. VIGLEIK ANGELTVEIT. TOPICS IN ALGEBRAIC TOPOLOGY. Weeks five through eight.

5. LACI BABAI. APPRENTICE PROGRAM: NUMBER THEORY AND LIN-EAR ALGEBRA.